OpenStudy (anonymous):

@Miracrown Could you please check my answer? I got 9 .

OpenStudy (anonymous):

I used the cross multiply method to get 9 .

sammixboo (sammixboo):

Hmmm... You gotta multiply 1/5 by 225 to get your answer, I believe \(\rm \color{#20bd23}{\dfrac{1}{5}\times225}\) Or you can do this \(\rm \color{#20bd23}{\dfrac{1\times225}{5}}\)

sammixboo (sammixboo):

You are trying to find 1/5 of 225

OpenStudy (anonymous):

225/5 = 9 ....

sammixboo (sammixboo):

Try again :)

sammixboo (sammixboo):

Because 9 x 5 = 45, not 223

sammixboo (sammixboo):

225*

Miracrown (miracrown):